# A method to derive self-consistent NLTE astrophysical parameters for 4 million high-resolution 4MOST stellar spectra in half a day with invertible neural networks

### v1

@@ Abstract
Context. Modern spectroscopic surveys obtain spectra for millions of stars. However, classical spectroscopic methods can often be computationally expensive, rendering them impractical for the analysis of large datasets.
Aims. We introduce a novel simulation-based deep-learning approach for the efficient analysis of high-resolution stellar spectra to be obtained with the upcoming high-resolution 4MOST spectrograph.
Methods. We used a suite of synthetic non-local thermodynamic equilibrium (NLTE) spectra generated with Turbospectrum to mimic 4MOST observations and trained a conditional invertible neural network (cINN) for the purpose of predicting self-consistently stellar surface parameters and chemical abundances. The cINN is a neural network architecture that estimates full posterior distributions for the target stellar properties, providing an intrinsic uncertainty estimate. We evaluated the predictive performance of the trained cINN model on both synthetic data and observed spectra of stars.
Results. We found that our new cINN trained on NLTE synthetic spectra is capable of recovering stellar parameters with average errors ($\sigma$) of $33$ K for $T_{\rm eff}$, $0.16$ dex for $\log(g)$, and $0.12$ dex for [Fe/H], $0.1$ dex for [Ca/Fe], $0.11$ for [Mg/Fe], and $0.51$ dex for [Li/Fe], respectively, at a signal to noise ratio of 250 per Angstrom. From the analysis of the observed spectra of Gaia-ESO / 4MOST / PLATO benchmark stars, we verified that our NLTE estimates for stellar parameters and abundances are consistent with results obtained with the independent code TSFitPy. We conclude that the NLTE cINN is robust and can, theoretically, evaluate 4 million high-resolution 4MOST spectra in less than a day, using GPU acceleration.
@@
stars:atmospheres – stars: abundances – techniques: spectroscopic – methods:statistical
@@ 1 Introduction
The past decade has seen revolutionary advances in observational stellar and Galactic astronomy. Large-scale stellar spectroscopic surveys, such as Gaia (Gaia Collaboration et al., 2016) and Gaia-ESO (Gilmore et al., 2012, 2022; Randich et al., 2013, 2022), GALAH (De Silva et al., 2015), LAMOST (Zhao et al., 2012), and APOGEE (Majewski et al., 2016, 2017) have provided millions of stellar spectra, and upcoming facilities like WEAVE (Jin et al., 2024), 4MOST (4-metre Multi-Object Spectroscopic Telescope; de Jong et al., 2019), and SDSS-V (Kollmeier et al., 2026) will further increase the data quantity by several orders of magnitude.
@@
To make use of this enormous wealth in observational data, all of these spectroscopic datasets require efficient analysis approaches. Classic fitting tools such as MOOG (Sneden, 1973), SIU (Reetz, 1991), PySME (Wehrhahn et al., 2023), KORG (Wheeler et al., 2023), and TSFitPy (Gerber et al., 2023; Storm and Bergemann, 2023) are highly flexible in terms of micro-physics. However, these codes typically require hours of CPU time in order to provide a detailed analysis of chemical abundances and stellar parameters for a single star. Although this approach is highly desirable and justified for individual stars, it becomes prohibitively expensive in terms of computation time when homogeneous, precise, and fast analysis of millions of high-resolution stellar spectra is required.
@@
Consequently, approaches that employ machine learning have been gaining considerable traction in recent years as a more efficient alternative. In particular, neural network based methods have proven to be promising. Ting et al. (2019) introduced ”the Payne”, which is a spectral fitting tool that uses a simple feed-forward neural network to facilitate a precise interpolation of synthetic spectra in a very high-dimensional parameter space, and was applied, for instance, in the analysis of GALAH DR4 (Buder et al., 2025). Another data-driven spectral analysis approach is known as ”the Cannon” (Ness et al., 2015), which uses real, well analysed spectra to build an efficient generative model for the probability distribution of the flux per wavelength given the stellar parameters, and was adapted for APOGEE spectra in Ness et al. (2016). The application of convolutional neural networks (CNN) has also been explored in, for instance, the analysis of RAVE spectra (Guiglion et al., 2020) and Gaia RVS spectra (Guiglion et al., 2024). The analysis of spectra has also seen exploration of various generative deep learning approaches. Liu et al. (2024) introduced the StellarGAN, a spectral classification approach for O to M-type stars based on generative adversarial networks (GANs; Goodfellow et al., 2014) that is trained on SDSS and APOGEE data. Another GAN-based approach is the GANDALF framework (Santoveña et al., 2024; Manteiga et al., 2025), which is trained on APOGEE and GALAH spectra and employs an autoencoder to embed spectra in a lower-dimensional representation space prior to deriving stellar atmospheric properties. More recently, Cvrček et al. (2025) presented a method that uses variational autoencoders (VAEs; Kingma and Welling, 2013; Rezende et al., 2014) to both predict stellar parameters and simulate spectra, which they trained on both real and synthetic spectra from the HARPS (Mayor et al., 2003) instrument. Gebran and Bentley (2025) also employed conditional VAEs (cVAEs; Sohn et al., 2015) to build a physics-aware surrogate model for spectral synthesis, trained on synthetic spectra from SYNSPEC (Hubeny and Lanz, 2011). However, most of these approaches rely on stellar parameters and chemical abundances, obtained using highly-simplified stellar synthetic spectra computed under assumptions of local thermodynamic equilibrium (LTE). As demonstrated in numerous recent studies (e.g. Korn et al., 2003; Asplund, 2005; Caffau et al., 2010; Lind et al., 2011; Bergemann et al., 2012b, a; Asplund et al., 2021; Amarsi et al., 2020; Lind and Amarsi, 2024; Bergemann and Hoppe, 2025), derivation of precise and accurate stellar parameters and abundance requires that stellar spectroscopic model grids take into account the effects of non-local thermodynamic equilibrium (NLTE).
@@
Here we present a novel approach to provide accurate astrophysical characterisation for gigantic sets of data, specifically from the upcoming 4MOST high-resolution disc and bulge survey 4MIDABLE-HR (Bensby et al., 2019). In this survey (PI: Bensby T. and Bergemann M.), we will obtain around four million spectra at high-resolution ($\mathrm{R}=20\,000$) in order to study the evolution of the Galactic disc and bulge. Dictated by the need to have highly efficient, stable, and adaptable algorithms, as well as the need to include NLTE in the chemical analysis (Lind and Amarsi, 2024; Bergemann and Hoppe, 2025), we are expanding upon the code that is presented in Kovalev et al. (2019) and applied to the Gaia-ESO survey therein. In particular, we introduce a novel simulation-based inference approach that for the first time combines NLTE synthetic spectra with a conditional invertible neural network (cINN; Ardizzone et al., 2019a, b). cINNs are a deep learning architecture that belongs to the generative normalising flow approaches and are specifically tailored towards solving inverse problems. They estimate the full posterior distribution, providing both an intrinsic uncertainty measure for the predictions and highlighting potential degeneracies in the given inverse problem. Although VAEs and GANs can be employed in a similar fashion (as in e.g. Cvrček et al., 2025), INNs have been shown to return better calibrated posterior distributions (Ardizzone et al., 2019a) and to be more robust to mode collapse (Ardizzone et al., 2019b; Kobyzev et al., 2021; Bond-Taylor et al., 2022). In addition, cINNs have already seen a series of successful applications in astronomy. These include the characterisation of stars from photometry (Ksoll et al., 2020), the analysis of emission lines from HII regions (Kang et al., 2022, 2023a), the inference of cosmic ray origins (Bister et al., 2022), the evaluation of exoplanet observations (Haldemann et al., 2023), the prediction of galaxy merger histories (Eisert et al., 2023) and the 3D reconstruction of dust distributions in star-forming cores (Ksoll et al., 2024). More relevant to the work presented in this study, cINNs have also already been proven to work well for the analysis of stellar spectra in several applications. This includes the evaluation of MUSE spectra based on LTE stellar atmosphere models from the PHOENIX library (Kang et al., 2023b, 2025) and the characterisation of high resolution spectra using a cINN trained on observed spectra of Milky Way field and cluster stars obtained with the GIRAFFE spectrograph in the Gaia-ESO survey (Candebat et al., 2024).
@@
In the following, we will outline our development of the first self-consistent cINN method for the analysis of observations of stars with the high-resolution spectrograph of 4MOST that is based fully on 25-D NLTE stellar spectral model grids. In Sect. 2, we introduce the training data for our cINN model, as well as a set of real benchmark spectra that we use to ascertain the performance of our model. Section 3 outlines the basic concepts of the cINN approach and implementation details for the specific inverse problem that we aim to solve in this study. In Sect. 4, we then describe the performance of our approach on both synthetic test data, as well as on archival spectra of the Gaia-ESO/4MOST/PLATO benchmark stars. Lastly, Sect. 5 provides a summary of our main results and outlook for the subsequent development of our method.
@@ 2 Data
In this section, we briefly outline the preparation of the training data (Sect. 2.1) and the composition of the set of real benchmark spectra used in Sect. 4.3 to gauge the performance of our trained model on real data (Sect. 2.2).
@@ 2.1 Training data
Figure 1: Prior distributions of the target parameters in the training data. The panels in the top two rows show 2D histograms of the correlation between selected stellar properties. The panels in the bottom three rows show 1D histograms of all target parameters.
@@ 2.2 Benchmark data
As 4MOST is not yet operational, we cannot benchmark our approach on real 4MOST data. To get an idea of how the cINN performs on observed spectra, we, therefore, used archival spectra of selected well-studied benchmark stars in addition to some other supplementary bright targets from our previous works. The spectra of the Gaia-ESO benchmark stars (Jofré et al., 2014; Heiter et al., 2015) were taken from Blanco-Cuaresma et al. (2014). These stars were observed with different astronomical facilities, including the Ultraviolet and Visual Echelle Spectrograph (UVES) at the 8m Very Large Telescope (VLT) of the European Southern Observatory (ESO) in Chile and the Narval spectrograph at the 2.03m Telescope Bernard Lyot (TBL) on the Pic du Midi de Bigorre in France. We note that some of the UVES spectra stem from the UVES Paranal Observatory Project (UVES, Bagnulo et al., 2003). Other stars were analysed in Fuhrmann et al. (1993); Fuhrmann (1998); Gehren et al. (2004, 2006); Bergemann and Gehren (2008); Bergemann and Cescutti (2010); Bergemann (2011), based on the spectroscopic data collected with the fibre optics Cassegrain échelle spectrograph (FOCES) at the 2.2m Calar-Alto telescope. Table 1 lists the set of sources that we selected together with literature results for $T_{\rm eff}$, $\log(g)$ and [Fe/H] (with the corresponding references). In total, our test set consisted of 58 unique stars across 72 observed spectra, including 39 main sequences stars, 11 subgiants, and 8 giants from K to F-type. All spectra were continuum normalised and corrected for radial velocity.
@@
All of the benchmark spectra have a higher native resolution than 4MOST, exist on a different wavelength grid than the output of the 4MOST high resolution spectrograph, and have varying S/N ratios. To process them with the cINN approach built in this work, we had to first 4MOST-ify the spectra. This included a reduction of the resolution to $\mathrm{R}=20\,000$ (matching the target resolution of the 4MIDABLE-HR survey) via convolution with an appropriate Gaussian kernel and rebinning the spectra to the wavelength bin grid of the high resolution 4MOST spectrograph. As most of the spectra in our test set do not extend far enough into the blue to cover the blue arm ($3926$ to $4355$ Å) of the 4MOST high resolution spectrograph, the latter rebinning step was limited to the green ($5160$ to $5730$ Å) and red arm ($6100$ to $6790$ Å) only. Consequently, the network presented in this study, operates only on these two wavelength windows. While the cINN described further below was trained to handle different S/N ratios in the input data, it is limited to the discrete set of S/N values currently generated in the training data (see Sect. 2.1). For the purposes of the benchmark conducted in Sect. 4.3, we, therefore, decided to degrade all spectra to a fixed $(\mathrm{S/N})_{\mathrm{fixed}}=250\,$Å-1. All spectra considered have a native S/N ratio that is higher, so that degrading the S/N (per wavelength bin) to that value was straight forward via the addition of Gaussian noise following

### v2

# A method to derive self-consistent NLTE astrophysical parameters for four million high-resolution 4MOST stellar spectra in half a day with invertible neural networks

@@ Abstract
Context. Modern spectroscopic surveys have the capacity to obtain the spectra of millions of stars. However, classical spectroscopic methods can often be computationally expensive, rendering them impractical for the analysis of large datasets.
Aims. We introduce a novel simulation-based, deep-learning approach for the efficient analysis of high-resolution stellar spectra that will be obtained with the upcoming high-resolution 4MOST spectrograph.
Methods. We used a suite of synthetic non-local thermodynamic equilibrium (NLTE) spectra generated with Turbospectrum to mimic 4MOST observations and trained a conditional invertible neural network (cINN) for the purpose of predicting self-consistently stellar surface parameters and chemical abundances. The cINN is a neural network architecture that estimates full posterior distributions for the target stellar properties, providing an intrinsic uncertainty estimate. We evaluated the predictive performance of the trained cINN model on both synthetic data and the observed spectra of stars.
Results. We found that our new cINN trained on NLTE synthetic spectra is capable of recovering stellar parameters with average errors ($\sigma$) of $33$ K for $T_{\rm eff}$, $0.16$ dex for $\log(g)$, and $0.12$ dex for [Fe/H], $0.1$ dex for [Ca/Fe], $0.11$ for [Mg/Fe], and $0.51$ dex for [Li/Fe], respectively, at a signal-to-noise ratio (S/N) of 250 per Angstrom. From the analysis of the observed spectra of Gaia-ESO / 4MOST / PLATO benchmark stars, we verified that our NLTE estimates for stellar parameters and abundances are consistent with results obtained with the independent code TSFitPy. We conclude that the NLTE cINN is robust and that it can, in theory, evaluate four million high-resolution 4MOST spectra in less than a day, using GPU acceleration.
@@
methods:statistical – techniques: spectroscopic – stars: abundances – stars: atmospheres
@@ 1 Introduction
The past decade has seen revolutionary advances in observational stellar and Galactic astronomy. Large-scale stellar spectroscopic surveys, such as Gaia (Gaia Collaboration et al., 2016) and Gaia-ESO (Gilmore et al., 2012, 2022; Randich et al., 2013, 2022), GALAH (De Silva et al., 2015), LAMOST (Zhao et al., 2012), and APOGEE (Majewski et al., 2016, 2017) have provided millions of stellar spectra, and upcoming facilities such as WEAVE (Jin et al., 2024), 4MOST (4-metre Multi-Object Spectroscopic Telescope; de Jong et al., 2019), and SDSS-V (Kollmeier et al., 2026) will further increase the data quantity by several orders of magnitude.
@@
To make use of this enormous wealth in observational data, all of these spectroscopic datasets require efficient analysis approaches. Classic fitting tools such as MOOG (Sneden, 1973), SIU (Reetz, 1991), PySME (Wehrhahn et al., 2023), KORG (Wheeler et al., 2023), and TSFitPy (Gerber et al., 2023; Storm and Bergemann, 2023) are highly flexible in terms of micro-physics. However, these codes typically require hours of CPU time in order to provide a detailed analysis of chemical abundances and stellar parameters for a single star. Although this approach is highly desirable and justified for individual stars, it becomes prohibitively expensive in terms of computation time when homogeneous, precise, and fast analyses of millions of high-resolution stellar spectra are required.
@@
Consequently, approaches that employ machine learning have been gaining considerable traction in recent years as a more efficient alternative. In particular, neural network-based methods have proven to be promising. Ting et al. (2019) introduced a spectral fitting tool known as ’the Payne’, which uses a simple feed-forward neural network to facilitate a precise interpolation of synthetic spectra in a very high-dimensional parameter space. It was applied, for instance, in the analysis of GALAH DR4 (Buder et al., 2025). Another data-driven spectral analysis approach is known as ’the Cannon’ (Ness et al., 2015), which uses real, well-analysed spectra to build an efficient generative model for the probability distribution of the flux per wavelength given the stellar parameters. It was adapted for APOGEE spectra in Ness et al. (2016). The application of convolutional neural networks (CNNs) has also been explored in, for instance, the analysis of RAVE spectra (Guiglion et al., 2020) and Gaia RVS spectra (Guiglion et al., 2024). The analysis of spectra has also seen exploration of various generative deep learning approaches. Liu et al. (2024) introduced the StellarGAN, a spectral classification approach for O to M-type stars based on generative adversarial networks (GANs; Goodfellow et al., 2014) that is trained on SDSS and APOGEE data. Another GAN-based approach is the GANDALF framework (Santoveña et al., 2024; Manteiga et al., 2025), which is trained on APOGEE and GALAH spectra and employs an autoencoder to embed spectra in a lower dimensional representation space prior to deriving stellar atmospheric properties. More recently, Cvrček et al. (2025) presented a method that uses variational autoencoders (VAEs; Kingma and Welling, 2013; Rezende et al., 2014) to both predict stellar parameters and simulate spectra, which they trained on both real and synthetic spectra from the HARPS (Mayor et al., 2003) instrument. Gebran and Bentley (2025) also employed conditional VAEs (cVAEs; Sohn et al., 2015) to build a physics-aware surrogate model for spectral synthesis, trained on synthetic spectra from SYNSPEC (Hubeny and Lanz, 2011). However, most of these approaches rely on stellar parameters and chemical abundances, obtained using highly simplified stellar synthetic spectra computed under assumptions of local thermodynamic equilibrium (LTE). As demonstrated in numerous recent studies (e.g. Korn et al., 2003; Asplund, 2005; Caffau et al., 2010; Lind et al., 2011; Bergemann et al., 2012b, a; Asplund et al., 2021; Amarsi et al., 2020; Lind and Amarsi, 2024; Bergemann and Hoppe, 2025), derivation of precise and accurate stellar parameters and abundance requires that stellar spectroscopic model grids take into account the effects of non-LTE (NLTE).
@@
Here, we present a novel approach to provide accurate astrophysical characterisation for gigantic sets of data; specifically, from the upcoming 4MOST high-resolution disc and bulge survey 4MIDABLE-HR (Bensby et al., 2019). In this survey (PI: Bensby T. and Bergemann M.), around four million spectra will be obtained at high resolution ($\mathrm{R}=20\,000$) to study the evolution of the Galactic disc and bulge. Dictated by the need to have highly efficient, stable, and adaptable algorithms, as well as the need to include NLTE in the chemical analysis (Lind and Amarsi, 2024; Bergemann and Hoppe, 2025), our aim is to expand upon the code presented in Kovalev et al. (2019) and applied to the Gaia-ESO survey therein. In particular, we introduced a novel simulation-based inference approach that for the first time combines NLTE synthetic spectra with a conditional invertible neural network (cINN; Ardizzone et al., 2019a, b). Overall, cINNs are a deep learning architecture that belongs to the generative normalising flow approaches and are specifically tailored towards solving inverse problems. They estimate the full posterior distribution, providing both an intrinsic uncertainty measure for the predictions and highlighting potential degeneracies in the given inverse problem. Although VAEs and GANs can be employed in a similar fashion (as in e.g. Cvrček et al., 2025), INNs have been shown to return better calibrated posterior distributions (Ardizzone et al., 2019a) and to be more robust to mode collapse (Ardizzone et al., 2019b; Kobyzev et al., 2021; Bond-Taylor et al., 2022). In addition, cINNs have already seen a series of successful applications in astronomy. These include the characterisation of stars from photometry (Ksoll et al., 2020), analysis of emission lines from HII regions (Kang et al., 2022, 2023a), inference of cosmic ray origins (Bister et al., 2022), evaluation of exoplanet observations (Haldemann et al., 2023), prediction of galaxy merger histories (Eisert et al., 2023), and 3D reconstruction of dust distributions in star-forming cores (Ksoll et al., 2024). In the role that is most relevant to the work presented in this study, cINNs have been proven to work well for the analysis of stellar spectra in several applications. This includes the evaluation of MUSE spectra based on LTE stellar atmosphere models from the PHOENIX library (Kang et al., 2023b, 2025) and the characterisation of high resolution spectra using a cINN trained on observed spectra of Milky Way field and cluster stars obtained with the GIRAFFE spectrograph in the Gaia-ESO survey (Candebat et al., 2024).
@@
In the following, we outline our development of the first self-consistent cINN method for the analysis of observations of stars with the high-resolution spectrograph of 4MOST that is based fully on 25-D NLTE stellar spectral model grids. In Sect. 2, we introduce the training data for our cINN model, as well as a set of real benchmark spectra that we use to ascertain the performance of our model. Section 3 outlines the basic concepts of the cINN approach and implementation details for the specific inverse problem that we aim to solve in this study. In Sect. 4, we then describe the performance of our approach on both synthetic test data, as well as on archival spectra of the Gaia-ESO/4MOST/PLATO benchmark stars. Lastly, Sect. 5 provides a summary of our main results and outlook for the subsequent development of our method.
@@ 2 Data
In this section, we briefly outline the preparation of the training data (Sect. 2.1). We also describe the composition of the set of real benchmark spectra used in Sect. 4.3 to gauge the performance of our trained model on real data (Sect. 2.2).
@@ 2.1 Training data
Figure 1: Prior distributions of the target parameters in the training data. The panels in the top two rows show 2D histograms of the correlation between selected stellar properties. The panels in the bottom three rows show the 1D histograms of all target parameters.
@@ 2.2 Benchmark data
As 4MOST is not yet operational, we cannot benchmark our approach on real 4MOST data. To get an idea of how the cINN performs on observed spectra, we therefore used archival spectra of selected well-studied benchmark stars in addition to some other supplementary bright targets from our previous works. The spectra of the Gaia-ESO benchmark stars (Jofré et al., 2014; Heiter et al., 2015) were taken from Blanco-Cuaresma et al. (2014). These stars were observed with different astronomical facilities, including the Ultraviolet and Visual Echelle Spectrograph (UVES) at the 8m Very Large Telescope (VLT) of the European Southern Observatory (ESO) in Chile and the Narval spectrograph at the 2.03m Telescope Bernard Lyot (TBL) on the Pic du Midi de Bigorre in France. We note that some of the UVES spectra stem from the UVES Paranal Observatory Project (UVES POP, Bagnulo et al., 2003). Other stars were analysed in Fuhrmann et al. (1993); Fuhrmann (1998); Gehren et al. (2004, 2006); Bergemann and Gehren (2008); Bergemann and Cescutti (2010); Bergemann (2011), based on the spectroscopic data collected with the fibre optics Cassegrain échelle spectrograph (FOCES) at the 2.2m Calar-Alto telescope. Table 3 in the appendix lists the set of sources that we selected together with literature results for $T_{\rm eff}$, $\log(g)$ and [Fe/H] (with the corresponding references). In total, our test set consisted of 58 unique stars across 72 observed spectra, including 39 main sequences stars, 11 subgiants, and 8 giants from the K to F types. All the spectra were continuum normalised and corrected for radial velocity.
@@
All of the benchmark spectra have a higher native resolution than 4MOST, existing on a different wavelength grid than the output of the 4MOST high resolution spectrograph and have varying S/Ns. To process them with the cINN approach built in this work, we had to first 4MOST-ify the spectra. This included a reduction of the resolution to $\mathrm{R}=20\,000$ (matching the target resolution of the 4MIDABLE-HR survey) via convolution with an appropriate Gaussian kernel and rebinning the spectra to the wavelength bin grid of the high resolution 4MOST spectrograph. As most of the spectra in our test set do not extend far enough into the blue to cover the blue arm ($3926$ to $4355$ Å) of the 4MOST high resolution spectrograph, the latter rebinning step was limited to the green ($5160$ to $5730$ Å) and red ($6100$ to $6790$ Å) arms only. Consequently, the network presented in this study, operates only on these two wavelength windows. While the cINN described further below was trained to handle different S/Ns in the input data, it is limited to the discrete set of S/N values currently generated in the training data (see Sect. 2.1). For the purposes of the benchmark conducted in Sect. 4.3, we decided to degrade all spectra to a fixed $(\mathrm{S/N})_{\mathrm{fixed}}=250\,$Å-1. All spectra considered have a native S/N that is higher, so that degrading the S/N (per wavelength bin) to that value was straightforward via the addition of Gaussian noise following